OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1

OpenStudy (anonymous):

Through elimination eq1-eq2 =2x=2 =x=1

OpenStudy (anonymous):

huh

OpenStudy (decentnabeel):

Equation1:−2x + y = 1 Equation2:−4x + y = −1 subtarct Equation1-Equation2 2x=2 x=1 put the value of y in equation1 -2(1)+y=1 y=1+2 y=3

OpenStudy (decentnabeel):

any confusion @Naenae1999